OpenStudy (anonymous):

Consider f(x)=sqrtx. Find all value(s), c , in the interval [0,1] such that the slope of the tangent line to the graph of f at c is parallel to the secant line through the points (0,f(0)) and (1,f(1)),

OpenStudy (anonymous):

first we need the slope of the line between \((0,f(0))\) and \((1,f(1))\) we have \(f(x)=\sqrt{x}\) so \(f(0)=\sqrt{0}=0\) and \(f(1)=\sqrt{1}=1\) slope is therefore 1

OpenStudy (anonymous):

now we need the formula for the slope of the tangent line that is the derivative, and the derivative of \(f(x)=\sqrt{x}\) is \(f'(x)=\frac{1}{2\sqrt{x}}\)

OpenStudy (anonymous):

finally we need the value of \(c\) for which the derivative is equal to the slope of the secant line, so we set \[\frac{1}{2\sqrt{c}}=1\] and solve for \(c\) that should take at most two steps

OpenStudy (anonymous):

thanks a bunch!